# Orbis: Overcoming Challenges of Long-Horizon Prediction in Driving World Models

### v1

@@ D.1 Trajectory evaluation metrics
To evaluate the distributional fidelity of generated trajectories, we use two primary metrics: pointwise error and curve similarity. These metrics serve as distance measures to evaluate distributional fidelity and coverage using precision–recall [38] in the driving trajectory space relevant to world model evaluation. Specifically, we represent a driving trajectory as a sequence of extrinsic transformation matrices $\boldsymbol{\mathcal{T}}=(\mathbf{T}_{1},\dots,\mathbf{T}_{T})$, where each $\mathbf{T}_{t}$ comprises a rotation (orientation) $\mathbf{R}_{t}\in SO(3)$ and a translation (position) $\mathbf{p}_{t}\in\mathbb{R}^{3}$, arranged as $\mathbf{T}_{t}=[\mathbf{R}_{t},\mathbf{p}_{t};\mathbf{0},1]$. For computing the ADE and Fréchet distances, we consider only the planar positions $\mathbf{p}_{t}\in\mathbb{R}^{2}$. Other parameters within $\mathbf{T}_{t}$, such as the rotation $\mathbf{R}_{t}$, can additionally be utilized to assess realism aspects like turning behavior and orientation evolution over time.
@@
Precision and Recall in Trajectory Embedding Space. To evaluate the distributional alignment between real and generated trajectories, we utilize the precision–recall framework [38]. We first map videos into the trajectory space using an inverse dynamics model (VGGT [71]). Let =$\mathcal{R}=\{\mathbf{r}_{i}\}_{i=1}^{N}$ and $\mathcal{G}=\{\mathbf{g}_{j}\}_{j=1}^{M}$ denote the planar positions trajectories of real and generated trajectories, respectively. For each real trajectory $\mathbf{r}_{i}$, define the threshold $\delta_{i}^{\mathcal{R}}$ as the distance to its $k$-th nearest neighbor in the real trajectory space $\mathcal{R}\setminus\{\mathbf{r}_{i}\}$. Conversely, for each generated trajectory $\mathbf{g}_{i}$, define $\delta_{j}^{\mathcal{G}}$ as the distance to its $k$-th nearest neighbor in the generated trajectory space $\mathcal{G}\setminus\{\mathbf{g}_{j}\}$. Precision and recall for a distance metric $d(\cdot,\cdot)$ (e.g., Fréchet) are then defined as:
@@
D.2 FVD evaluation.
@@ Appendix E More results
Cosmos [1] 10 210.56 249.08 244.80
Vista [15] 10 289.95 351.42 353.27
GEM [23] 10 348.36 218.61 318.73
DW* [29] 5 298.97 N/A 334.89

### v2

@@ D.1 Trajectory evaluation metrics
To evaluate the distributional fidelity of generated trajectories, we use two primary metrics: pointwise error and curve similarity. These metrics serve as distance measures to evaluate distributional fidelity and coverage using precision–recall [kynkaanniemi2019pr] in the driving trajectory space relevant to world model evaluation. Specifically, we represent a driving trajectory as a sequence of extrinsic transformation matrices $\boldsymbol{\mathcal{T}}=(\mathbf{T}_{1},\dots,\mathbf{T}_{T})$, where each $\mathbf{T}_{t}$ comprises a rotation (orientation) $\mathbf{R}_{t}\in SO(3)$ and a translation (position) $\mathbf{p}_{t}\in\mathbb{R}^{3}$, arranged as $\mathbf{T}_{t}=[\mathbf{R}_{t},\mathbf{p}_{t};\mathbf{0},1]$. For computing the ADE and Fréchet distances, we consider only the planar positions $\mathbf{p}_{t}\in\mathbb{R}^{2}$. Other parameters within $\mathbf{T}_{t}$, such as the rotation $\mathbf{R}_{t}$, can additionally be utilized to assess realism aspects like turning behavior and orientation evolution over time.
@@
Precision and Recall in Trajectory Embedding Space. To evaluate the distributional alignment between real and generated trajectories, we utilize the precision–recall framework [kynkaanniemi2019pr]. We first map videos into the trajectory space using an inverse dynamics model (VGGT [wang2025vggt]). Let $\mathcal{R}=\{\mathbf{r}_{i}\}_{i=1}^{N}$ and $\mathcal{G}=\{\mathbf{g}_{j}\}_{j=1}^{M}$ denote the planar positions trajectories of real and generated trajectories, respectively. For each real trajectory $\mathbf{r}_{i}$, define the threshold $\delta_{i}^{\mathcal{R}}$ as the distance to its $k$-th nearest neighbor in the real trajectory space $\mathcal{R}\setminus\{\mathbf{r}_{i}\}$. Conversely, for each generated trajectory $\mathbf{g}_{i}$, define $\delta_{j}^{\mathcal{G}}$ as the distance to its $k$-th nearest neighbor in the generated trajectory space $\mathcal{G}\setminus\{\mathbf{g}_{j}\}$. Precision and recall for a distance metric $d(\cdot,\cdot)$ (e.g., Fréchet) are then defined as:
@@
D.2 VQA evaluation metrics.
PSNR and SSIM.
As shown in the Table 10, we computed average PSNR and SSIM metrics over two shifted video windows of 10 frames (i.e. 2 seconds at 5fps) over 400 generated videos on nuPlan-turns evaluation benchmark. We resize generated videos of all methods to same resolution for a fair comparison. We observe that Orbis model performs marginally better than other methods over the first window. However, all methods converge to similarly low numbers in the second window.
Table 10: VQA metrics: PSNR and SSIM on windowed video of 10 frames over 400 generated videos on nuPlan-turns.
PSNR SSIM
Model frames 0-9 frames 10-19 frames 0-9 frames 10-19
Cosmos 17.29 13.38 0.47 0.38
GEM 14.85 13.73 0.42 0.41
Vista 15.04 13.70 0.44 0.42
DW 17.67 14.70 0.44 0.38
Orbis 18.72 14.75 0.52 0.42
DOVER.
We evaluated the models on the DOVER [wu2023exploring] score, which is also used in the data curation/filtering pipeline of Cosmos. In the Table 11, we report a comparison of the DOVER scores on the nuPlan Turns dataset, computed over 17s long videos. We include both the results for the full Cosmos pipeline (including the extra 7B-parameters diffusion refiner, "Cosmos+ref"), and for the pure Cosmos world model with a non-generative decoder.
Table 11: Blind VQA DOVER metric on 400 generated videos of 17s on nuPlan-turns.
[…]
Cosmos 19.94
Cosmos+refinement 28.99
GEM 19.76
Vista 21.14
DW 21.92
Orbis 21.34
D.3 FVD evaluation.
@@ Appendix E More results
Cosmos [agarwal2025cosmos] 10 210.56 249.08 244.80
Vista [gao2024vista] 10 289.95 351.42 353.27
GEM [hassan2024gem] 10 348.36 218.61 318.73
DW* [hu2024drivingworld] 5 298.97 N/A 334.89
@@
Inference speed and memory requirements.
Table 13 compares the average times (fps) needed to generate a frame, and the required GPU memory. Orbis has the best throughput compared to all other methods. This advantage can be attributed to the smaller size of the Orbis model. By parallelizing Orbis’ computation in batches we can achieve an higher throughput. GEM and Vista are based on the same architecture but use different sampling protocols trading off FPS and VRAM.
We also compare the inference speed of our discrete Orbis-MG model and the continuous Orbis-FM model. We also report the GPU (VRAM) memory requirements for both methods during inference. Orbis-MG shows better inference speed compared to the Orbis-FM model. However, since Orbis-FM achieves significantly better video generation performance, it remains the default choice despite the speed advantage of the discrete model.
Table 13: Comparison of inference speed and VRAM memory requirements of different models.
[…]
Cosmos 0.18 29
Vista 0.58 86
GEM 0.44 45
DW 0.25 10
Orbis-FM (ours) 0.70 24
Orbis-MG (ours) 0.85 21